# Information Inequalities for Five Random Variables

### v1

@@ 8. The minimal set of inequalities
Experimental results reported in Section 5 and discussed in Section 6 provided the complete list of five-variable non-Shannon entropy inequalities implied by the existence of an $n$-copy for $n\leqslant 9$. Two families of non-Shannon inequalities, generalizing the ones found experimentally, were proven, in Theorem 11 and Theorem 16, respectively, to hold in every polymatroid with an $n$-copy. We conjecture that these families actually characterize those $5$-variable polymatroids that have an $n$-copy, so no further non-Shannon inequalities can be discovered by the version of the Maximum Entropy Method utilized in this paper.
In the $acbd\mkern 1.5mu]<0$ case the family of non-Shannon inequalities provided by Theorem 16 matches exactly the inequalities obtained experimentally for $n\leqslant 9$.
In the $abcd\mkern 1.5mu]<0$ case the family provided by Theorem 11 is parametrized by the downward closed subsets $S$ of the diagonal set $D_{n}\subset\mathbb{N}\times\mathbb{N}$. Not all of the generated inequalities correspond to facets of the cone $\mathcal{Q}^{*}_{1}$. While they are valid non-Shannon inequalities, some of them are consequences of others. Table 4 shows the downward closed subsets of $D_{3}$ as well as the corresponding $\mathbf{v}_{S}=\langle\alpha_{S},\beta_{S},\gamma_{S}\rangle$ triplets from Definition 4. Two triplets, marked by *, are not in Table 3. The corresponding inequality
$(a,b\mkern 1.0mu{|}\mkern 1.0muz)+\alpha\big([abcd\mkern 1.5mu]+\mathcal{Z}\big)+\beta\mkern 1.5mu\mathcal{C}+\gamma\mathcal{D}\geqslant 0$ (27)
with $\alpha=5$ and $\beta=\gamma=3$ is the average of the inequalities obtained from the triplets numbered $6$, $10$ and $13$, thus it is a consequence of them. The main goal of this Section is to obtain a description of those downward closed subsets of $D_{n}$ that generate facets of $\mathcal{Q}^{*}_{1}$, that is, inequalities that are not consequences of the others.
$\scriptstyle 1$$\scriptstyle 2$$\scriptstyle 3$$\scriptstyle 4$$\scriptstyle 5$$\scriptstyle 6$$\scriptstyle 7$$\scriptstyle 8$$\scriptstyle 9$$\scriptstyle 10$$\scriptstyle 11$$\scriptstyle 12$$\scriptstyle 13$
1 $\langle 1,0,0\rangle$
2 $\langle 2,1,0\rangle$
3 $\langle 3,3,0\rangle$
4 $\langle 2,0,1\rangle$
5 $\langle 3,1,1\rangle$
6 $\langle 4,3,1\rangle$
7*7 $\langle 5,3,3\rangle$
8 $\langle 6,5,3\rangle$
9 $\langle 3,0,3\rangle$
10 $\langle 4,1,3\rangle$
11*11 $\langle 5,3,3\rangle$
12 $\langle 6,3,5\rangle$
13 $\langle 7,5,5\rangle$
Table 4. Downward closed subsets of $D_{3}$
Since the inequality (27) contains the fixed term $(a,b\mkern 1.0mu{|}\mkern 1.0muz)$, and trivially holds true when $abcd\mkern 1.5mu]+\mathcal{Z}\geqslant 0$, it is a consequence of the inequalities obtained from the triplets $\{\langle\alpha_{i},\beta_{i},\gamma_{i}\rangle:i\in I\}$ if there is a convex combination
@@ 13 $\langle 7,5,5\rangle$
Let $i_{1}<i_{2}<i_{3}$, and $j_{1}>j_{2}>j_{3}$. (i) $\mathbf{v}_{S}$ is superseded by the vectors $\{\mathbf{v}_{S+(i_{1},j_{1})},\mathbf{v}_{S-(i_{2},j_{2})},\mathbf{v}_{S+(i_{3},j_{3})}\}$ if and only if
@@
(ii) $\mathbf{v}_{S}$ is superseded by $\{\mathbf{v}_{S-(i_{1},j_{1})},\mathbf{v}_{S+(i_{2},j_{2})},\mathbf{v}_{S-(i_{3},j_{3})}\}$ if and only if
@@
We prove (i) only, (ii) is similar. Let $\mathsfit b\mkern 1.0mu_{1}=\mathsfit b\mkern 1.0mu(i_{1},j_{1})$, $\mathsfit b\mkern 1.0mu_{2}=\mathsfit b\mkern 1.0mu(i_{2},j_{2})$ and $\mathsfit b\mkern 1.0mu_{3}=\mathsfit b\mkern 1.0mu(i_{3},j_{3})$. Then, according to Definition 4,
$\mathbf{v}_{S+(i_{1},j_{1})}$ $=\langle\alpha_{S}+\mathsfit b\mkern 1.0mu_{1},\beta_{S}+i_{1}\mathsfit b\mkern 1.0mu_{1},\gamma_{S}+j_{1}\mathsfit b\mkern 1.0mu_{1}\rangle,$
$\mathbf{v}_{S-(i_{2},j_{2})}$ $=\langle\alpha_{S}-\mathsfit b\mkern 1.0mu_{2},\beta_{S}-i_{2}\mathsfit b\mkern 1.0mu_{2},\gamma_{S}-j_{2}\mathsfit b\mkern 1.0mu_{2}\rangle,$
$\mathbf{v}_{S+(i_{3},j_{3})}$ $=\langle\alpha_{S}+\mathsfit b\mkern 1.0mu_{3},\beta_{S}+i_{3}\mathsfit b\mkern 1.0mu_{3},\gamma_{S}+j_{3}\mathsfit b\mkern 1.0mu_{3}\rangle.$
@@
$\alpha_{S}$ $\leqslant\lambda_{1}(\alpha_{S}+\mathsfit b\mkern 1.0mu_{1})+\lambda_{2}(\alpha_{S}-\mathsfit b\mkern 1.0mu_{2})+\lambda_{3}(\alpha_{S}+\mathsfit b\mkern 1.0mu_{3}),$
$\beta_{S}$ $\geqslant\lambda_{1}(\beta_{S}+i_{1}\mathsfit b\mkern 1.0mu_{1})+\lambda_{2}(\beta_{S}-i_{2}\mathsfit b\mkern 1.0mu_{2})+\lambda_{3}(\beta_{S}+i_{3}\mathsfit b\mkern 1.0mu_{3}),$
$\gamma_{S}$ $\geqslant\lambda_{1}(\gamma_{S}+j_{1}\mathsfit b\mkern 1.0mu_{1})+\lambda_{2}(\gamma_{S}-j_{2}\mathsfit b\mkern 1.0mu_{2})+\lambda_{3}(\gamma_{S}+j_{3}\mathsfit b\mkern 1.0mu_{3}).$
@@
$\lambda_{2}\mathsfit b\mkern 1.0mu_{2}$ $\leqslant\lambda_{1}\mathsfit b\mkern 1.0mu_{1}+\lambda_{3}\mathsfit b\mkern 1.0mu_{3},$
$i_{2}\lambda_{2}\mathsfit b\mkern 1.0mu_{2}$ $\geqslant i_{1}\lambda_{1}\mathsfit b\mkern 1.0mu_{1}+i_{3}\lambda_{3}\mathsfit b\mkern 1.0mu_{3},$
$j_{2}\lambda_{2}\mathsfit b\mkern 1.0mu_{2}$ $\geqslant j_{1}\lambda_{1}\mathsfit b\mkern 1.0mu_{1}+j_{3}\lambda_{3}\mathsfit b\mkern 1.0mu_{3}.$
Clearly, $\lambda_{2}$ must be strictly positive as $\mathsfit b\mkern 1.0mu_{1}$, $\mathsfit b\mkern 1.0mu_{2}$, and $\mathsfit b\mkern 1.0mu_{3}$ are positive. Introducing $\mu_{1}=(\lambda_{1}\mathsfit b\mkern 1.0mu_{1})/(\lambda_{2}\mathsfit b\mkern 1.0mu_{2})$ and $\mu_{3}=(\lambda_{3}\mathsfit b\mkern 1.0mu_{3})/(\lambda_{2}\mathsfit b\mkern 1.0mu_{2})$, this system is equivalent to
@@
One can assume that the first inequality holds with an equality. Since $i_{1}<i_{2}<i_{3}$, the second inequality holds when $i_{2}$ is above the point which splits the interval $i_{1},i_{3}]$ in ratio $\mu_{3}$ to $\mu_{1}$. Similarly, $j_{1}<-j_{2}<-j_{3}$ implies that the third inequality holds when $j_{2}$ is below the point that splits $j_{1},-j_{3}]$ in the same ratio. Thus non-negative numbers $\mu_{1}$ and $\mu_{3}$ satisfying this three inequalities exist if and only if the proportion of $i_{2},i_{3}]$ in $i_{1},i_{3}]$ is not larger than the proportion of $j_{2},-j_{3}]$ in the interval $j_{1},-j_{3}]$, that is,
@@
A downward closed set $S\subset\mathbb{N}\times\mathbb{N}$ can be specified in two ways. Either by a non-increasing sequence $S^{\mathrm{col}}=(c_{0},c_{1},\dots,c_{k})$ specifying the maximal values in columns 0, …, $k$, or by a non-increasing sequence $S^{\mathrm{row}}=(r_{0},r_{1},\dots,r_{\ell})$ specifying the maximal values in rows 0, …, $\ell$. It is easy to see that
@@
If the vector $\mathbf{v}_{S}$ for some $S\subseteq D_{n}$ is not superseded by other vectors generated by subsets of $D_{n}$, then either the sequence $S^{\mathrm{col}}$ is strictly decreasing, or the sequence $S^{\mathrm{row}}$ is strictly decreasing.
@@
$\scriptstyle r$$\scriptstyle r_{1}$$\scriptstyle r_{2}$$\scriptstyle c_{1}$$\scriptstyle c_{2}$$\scriptstyle c$$\scriptstyle r$$\scriptstyle r_{1}$$\scriptstyle r_{2}$$\scriptstyle c$$\scriptstyle c_{1}$$\scriptstyle c_{2}$
Figure 1. Closest horizontal and vertical boundary segments
In the first case $c_{1}<c_{2}<c$, and $r_{1}<r_{2}<r$; in the second case $c<c_{1}<c_{2}$ and $r<r_{1}<r_{2}$. Apply Lemma 18 to the marked points, and observe that the modified downward closed sets are always subsets of $D_{n}$. In the first case
@@
By Corollary 20, the downward closed set corresponding to an extremal vertex is either a staircase with step heights $1$ (when $S^{\mathrm{row}}$ is strictly decreasing) which we call horizontal, or the mirror image of such a staircase. The only configuration that belongs to both cases is the diagonal $D_{n}$. It will be more convenient to use the column-sequence $(c_{0},c_{1},\dots,c_{k})$ to represent horizontal staircases. Here $k\geqslant 0$ is the length of the staircase, also denoted by $\operatorname{\mathrm{len}}(S)$. The last column size (height) is necessarily $c_{k}=0$, and $c_{i+1}$ equals either $c_{i}$ or $c_{i}\mkern 1.5mu{+}\mkern 1.5mu1$ for every $0\leqslant i<k$. In the rest of this section, all staircases, if not mentioned otherwise, are horizontal ones.
@@
The staircase $S$ is PNP-reducible in $D_{n}$ if there are $i_{1}<i_{2}<i_{3}$ and $j_{1}>j_{2}>j_{3}$ such that $S_{1}=S+(i_{1},j_{1})$, $S_{2}=S-(i_{2},j_{2})$, and $S_{3}=S+(i_{3},j_{3})$ are staircases in $D_{n}$ and $\mathbf{v}_{S}$ is superseded by $\{\mathbf{v}_{S_{1}},\mathbf{v}_{S_{2}},\mathbf{v}_{S_{3}}\}$. $S$ is PNP-irreducible if it is not PNP-reducible.
NPN-reducibility and NPN-irreducibility is defined analogously, using staircases $S-(i_{1},j_{1})$, $S+(i_{2},j_{2})$, and $S-(i_{3},j_{3})$, assuming that they are also subsets of $D_{n}$. $S$ is irreducible in $D_{n}$ if it is both PNP- and NPN-irreducible. Finally, let $\mathcal{S}_{n}$ be the collection of the irreducible staircases that are subsets of $D_{n}$.
@@
depending on whether the last to elements of the column sequence of $S$ are equal.
@@
(iv) If $S$ is reducible in $D_{n}$ while $S^{\prime}$ is not reducible in $D_{n-1}$, then the reduction must use $(i_{3},j_{3})$ which is in $D_{n}$ but not in $D_{n-1}$. If it is an NPN-reduction then it must use the newly added point $(n,0)$, in other cases the reduction can be shifted back to $S^{\prime}$. In the case of a PNP-reduction this additional point is (n1,1)(n-1,1) (when extending the staircase by a column of height zero), or can be shifted back to $S^{\prime}$ again. ∎
@@
Based on Claim 21, the incremental algorithm, sketched as Code 1, generates all horizontal irreducible staircases. The PNP- and NPN-irreducibility can be checked based on Lemma 18. The last point $(i_{3},j_{3})$ is fixed, and the naïve implementation requires quadratic running time in $\operatorname{\mathrm{len}}(S)$. With some simple bookkeeping it can be reduced to a backward scanning of the column sequence, resulting in linear running time.
Using the algorithm we have computed the complete set of irreducible staircases up to $n=60$. The number of new staircases that remained irreducible in each subsequent generation matches the sequence A103116 in the Encyclopedia of Integer Sequences [19]:
@@
where $\varphi$ is Euler’s totient function, which suggests that the connection is based on the number of different slopes determined by the lattice points in a rectangle. Proving the equivalence of these two sequences is an intriguing open problem.
For better visualization, triplets $\langle\alpha_{S},\beta_{S},\gamma_{S}\rangle$ corresponding to these irreducible staircases are plotted as the $3$-dimensional points $\langle\beta/\alpha,\gamma/\alpha,\alpha\rangle$ using logarithmic scale for the third $\alpha$ coordinate. The plot in Figure 2 contains all 126,981 extremal triplets in the range $\beta,\gamma\leqslant 20\alpha$. Some of the plotted triplets appear as late as generation $n=80$; later generations do not contribute to this part of the complete set. For comparison, some triplets in the $80$-th generation have values larger than $2^{85}$.
@@
Figure 2. Extremal configurations
@@
$\sum_{i\leqslant x,\;j\leqslant y}\mathsfit b\mkern 1.0mu(i,j)$ $\approx\mathsfit b\mkern 1.0mu(x\mkern 1.5mu{+}\mkern 1.5mu1,y\mkern 1.5mu{+}\mkern 1.5mu1),\mbox{ and}$
$\sum_{i\leqslant x,\;j\leqslant y}i\mkern 1.5mu\mathsfit b\mkern 1.0mu(i,j)$ $\approx x\mkern 1.5mu\mathsfit b\mkern 1.0mu(x\mkern 1.5mu{+}\mkern 1.5mu1,y\mkern 1.5mu{+}\mkern 1.5mu1),$
the vector $\mathbf{v}_{S(a,b)}$ is well approximated by $\mathsfit b\mkern 1.0mu(x\mkern 1.5mu{+}\mkern 1.5mu1,y\mkern 1.5mu{+}\mkern 1.5mu1)\langle 1,x,y\rangle$, where $(x,y)\in S(a,b)$ is the point where $\mathsfit b\mkern 1.0mu(x\mkern 1.5mu{+}\mkern 1.5mu1,y\mkern 1.5mu{+}\mkern 1.5mu1)$ takes its maximal value. As the function $\mathsfit b\mkern 1.0mu(x,y)$ strictly increases in both coordinates, this maximum is taken on the boundary diagonal of the right-angled triangle $S(a,b)$ that has endpoints $(0,b)$ and $(a,0)$. Using the Stirling formula $n!\approx\sqrt{2\pi n}\mkern 1.5mu(n/e)^{n}$, we have
$\mathsfit b\mkern 1.0mu(x,y)=\frac{\mkern 1.5mu(x+y)!\mkern 1.5mu}{x!\,y!}\approx\frac{1}{\sqrt{2\pi}}\frac{(x+y)^{x+y+1/2}}{x^{x+1/2}\mkern 1.5muy^{y+1/2}}.$
Introducing $\varphi(x)=(x+\frac{1}{2})\log x$, we see that the logarithm of $\mathsfit b\mkern 1.0mu(x,y)$ is well approximated by the function
@@
Figure 3. The $\theta(u+1,v+1)$ function

### v2

@@ 8. The Minimal Set of Inequalities
Experimental results reported in Section 5 and discussed in Section 6 provided the complete list of five-variable non-Shannon entropy inequalities implied by the existence of an $n$-copy for $n\leqslant 9$. Two families of non-Shannon inequalities, generalizing the ones found experimentally, were proven, in Theorem 11 and Theorem 16, respectively, to hold in every polymatroid with an $n$-copy. We conjecture that these families actually characterize those five-variable polymatroids that have an $n$-copy, so no further non-Shannon inequalities can be discovered by the version of the Maximum Entropy Method utilized in this paper.
In the $acbd\mkern 1.5mu]<0$ case, the family of non-Shannon inequalities provided by Theorem 16 matches exactly the inequalities obtained experimentally for $n\leqslant 9$.
In the $abcd\mkern 1.5mu]<0$ case, the family provided by Theorem 11 is parametrized by the downward closed subsets $s$ of the diagonal set $D_{n}\subset\mathbb{N}\times\mathbb{N}$. Unfortunately, not all of the obtained inequalities correspond to facets of the cone $\mathcal{Q}^{*}_{1}$. The reason is that, while all of them are valid entropy inequalities, some are consequences of the others. Recall that the inequality corresponding to the triplet $\langle\alpha,\beta,\gamma\rangle$ is
$(a,b\mkern 1.5mu{|}\mkern 1.5muz)+\alpha\big([abcd\mkern 1.5mu]+\mathcal{Z}\big)+\beta\mkern 1.5mu\mathcal{C}+\gamma\mathcal{D}\geqslant 0.$ (28)
As an example, the triplet $\langle 5,3,3\rangle$, obtained from the downward closed subsets numbered #7\#7 and #11\#11 in Table 5, does not appear in Table 3, the list of computational results. The corresponding inequality is indeed a consequence of those obtained from the triplets numbered #6\#6, #10\#10, and #13\#13. This is so, as $\langle 5,3,3\rangle$ is their average:
$\langle 5,3,3\rangle=\frac{1}{3}\big(\langle 4,3,1\rangle+\langle 4,1,3\rangle+\langle 7,5,5\rangle\big),$
and the same holds for the corresponding inequalities. The main goal of this section is to obtain a description of those downward closed subsets of $D_{n}$ that generate facets of $\mathcal{Q}^{*}_{1}$, that is, inequalities that are not consequences of the others.
Since the inequality (28) contains the fixed term $(a,b\mkern 1.5mu{|}\mkern 1.5muz)$, and trivially holds true when $abcd\mkern 1.5mu]+\mathcal{Z}\geqslant 0$, it is a consequence of the inequalities obtained from the triplets $\{\langle\alpha_{i},\beta_{i},\gamma_{i}\rangle:i\in I\}$ if and only if there is a convex linear combination
@@
Let $i_{1}<i_{2}<i_{3}$, and $j_{1}>j_{2}>j_{3}$.
  • (i)
    $\mathbf{v}_{S}$ is superseded by the vectors $\{\mathbf{v}_{S+(i_{1},j_{1})},\mathbf{v}_{S-(i_{2},j_{2})},\mathbf{v}_{S+(i_{3},j_{3})}\}$ if and only if
@@
  • (ii)
    $\mathbf{v}_{S}$ is superseded by $\{\mathbf{v}_{S-(i_{1},j_{1})},\mathbf{v}_{S+(i_{2},j_{2})},\mathbf{v}_{S-(i_{3},j_{3})}\}$ if and only if
@@
We prove (i) only, (ii) is similar. Let $\mathbfit b_{1}=\mathbfit b(i_{1},j_{1})$, $\mathbfit b_{2}=\mathbfit b(i_{2},j_{2})$ and $\mathbfit b_{3}=\mathbfit b(i_{3},j_{3})$. Then, according to Definition 4,
$\mathbf{v}_{S+(i_{1},j_{1})}$ $=\langle\alpha_{S}+\mathbfit b_{1},\beta_{S}+i_{1}\mathbfit b_{1},\gamma_{S}+j_{1}\mathbfit b_{1}\rangle,$
$\mathbf{v}_{S-(i_{2},j_{2})}$ $=\langle\alpha_{S}-\mathbfit b_{2},\beta_{S}-i_{2}\mathbfit b_{2},\gamma_{S}-j_{2}\mathbfit b_{2}\rangle,$
$\mathbf{v}_{S+(i_{3},j_{3})}$ $=\langle\alpha_{S}+\mathbfit b_{3},\beta_{S}+i_{3}\mathbfit b_{3},\gamma_{S}+j_{3}\mathbfit b_{3}\rangle.$
@@
$\alpha_{S}$ $\leqslant\lambda_{1}(\alpha_{S}+\mathbfit b_{1})+\lambda_{2}(\alpha_{S}-\mathbfit b_{2})+\lambda_{3}(\alpha_{S}+\mathbfit b_{3}),$
$\beta_{S}$ $\geqslant\lambda_{1}(\beta_{S}+i_{1}\mathbfit b_{1})+\lambda_{2}(\beta_{S}-i_{2}\mathbfit b_{2})+\lambda_{3}(\beta_{S}+i_{3}\mathbfit b_{3}),$
$\gamma_{S}$ $\geqslant\lambda_{1}(\gamma_{S}+j_{1}\mathbfit b_{1})+\lambda_{2}(\gamma_{S}-j_{2}\mathbfit b_{2})+\lambda_{3}(\gamma_{S}+j_{3}\mathbfit b_{3}).$
@@
$\lambda_{2}\mathbfit b_{2}$ $\leqslant\lambda_{1}\mathbfit b_{1}+\lambda_{3}\mathbfit b_{3},$
$i_{2}\lambda_{2}\mathbfit b_{2}$ $\geqslant i_{1}\lambda_{1}\mathbfit b_{1}+i_{3}\lambda_{3}\mathbfit b_{3},$
$j_{2}\lambda_{2}\mathbfit b_{2}$ $\geqslant j_{1}\lambda_{1}\mathbfit b_{1}+j_{3}\lambda_{3}\mathbfit b_{3}.$
Clearly, $\lambda_{2}$ must be strictly positive as $\mathbfit b_{1}$, $\mathbfit b_{2}$, and $\mathbfit b_{3}$ are positive. Introducing $\mu_{1}=(\lambda_{1}\mathbfit b_{1})/(\lambda_{2}\mathbfit b_{2})$ and $\mu_{3}=(\lambda_{3}\mathbfit b_{3})/(\lambda_{2}\mathbfit b_{2})$, this system is equivalent to
@@
One can assume that the first inequality holds with an equality. Since $i_{1}<i_{2}<i_{3}$, the second inequality holds when $i_{2}$ is above the point which splits the interval $i_{1},i_{3}]$ in ratio $\mu_{3}$ to $\mu_{1}$. Similarly, $j_{1}<-j_{2}<-j_{3}$ implies that the third inequality holds when $j_{2}$ is below the point that splits $j_{1},-j_{3}]$ in the same ratio. Thus, non-negative numbers $\mu_{1}$ and $\mu_{3}$ satisfying these three inequalities exist if and only if the proportion of $i_{2},i_{3}]$ in $i_{1},i_{3}]$ is not larger than the proportion of $j_{2},-j_{3}]$ in the interval $j_{1},-j_{3}]$, that is,
@@
A downward closed set $S\subset\mathbb{N}\times\mathbb{N}$ can be specified in two ways. Either by a non-increasing sequence $S^{\mathrm{col}}=(c_{0},c_{1},\dots,c_{k})$ specifying the maximal values in columns 0, …, $k$ (column sequence), or by a non-increasing sequence $S^{\mathrm{row}}=(r_{0},r_{1},\dots,r_{\ell})$ specifying the maximal values in rows 0, …, $\ell$ (row sequence). It is easy to see that
@@
For the downward closed subsets of $D_{3}$ the corresponding column and row sequences are presented in Table 6.
[…]
Table 6. Column and row sequences for subsets of $D_{3}$.
@@
If for some $S\subseteq D_{n}$ the vector $\mathbf{v}_{S}$ is not superseded by other vectors generated from subsets of $D_{n}$, then either the sequence $S^{\mathrm{col}}$ is strictly decreasing, or the sequence $S^{\mathrm{row}}$ is strictly decreasing.
@@
[…]
Figure 1. Nearest horizontal and vertical boundary segments of the gray region are marked by the arrows. Red circles indicate points to be added and subtracted.
In the first case $c_{1}<c_{2}<c$, and $r_{1}<r_{2}<r$; in the second case $c<c_{1}<c_{2}$ and $r<r_{1}<r_{2}$. Apply Lemma 18 to the marked points and observe that the modified downward closed sets are always subsets of $D_{n}$. In the first case
@@
By Corollary 20, the downward closed set corresponding to an extremal vertex is either a staircase with step heights $1$ (when $S^{\mathrm{row}}$ is strictly decreasing), which we call horizontal, or the mirror image of such a staircase. The only configuration that belongs to both cases is the diagonal $D_{n}$. It will be more convenient to use the column-sequence $(c_{0},c_{1},\dots,c_{k})$ to represent horizontal staircases. Here $k\geqslant 0$ is the length of the staircase, also denoted by $\operatorname{\mathrm{len}}(S)$. The last column size (height) is necessarily $c_{k}=0$, and $c_{i}$ equals either $c_{i+1}$ or $c_{i+1}\mkern 1.5mu{+}\mkern 1.5mu1$ for every $0\leqslant i<k$. In the rest of this section, all staircases, if not mentioned otherwise, are horizontal ones.
@@
The staircase $S$ is Positive-Negative-Positive (PNP)-reducible in $D_{n}$ if there are $i_{1}<i_{2}<i_{3}$ and $j_{1}>j_{2}>j_{3}$ such that $S_{1}=S+(i_{1},j_{1})$, $S_{2}=S-(i_{2},j_{2})$, and $S_{3}=S+(i_{3},j_{3})$ are staircases in $D_{n}$ and $\mathbf{v}_{S}$ is superseded by $\{\mathbf{v}_{S_{1}},\mathbf{v}_{S_{2}},\mathbf{v}_{S_{3}}\}$. $S$ is PNP-irreducible if it is not PNP-reducible.
Negative-Positive-Negative (NPN)-reducibility and NPN-irreducibility is defined analogously, using staircases $S-(i_{1},j_{1})$, $S+(i_{2},j_{2})$, and $S-(i_{3},j_{3})$, assuming that they are also subsets of $D_{n}$. $S$ is irreducible in $D_{n}$ if it is both PNP- and NPN-irreducible. Finally, let $\mathcal{S}_{n}$ be the collection of the irreducible staircases that are subsets of $D_{n}$.
@@
depending on whether the last two elements of the column sequence of $S$ are equal.
@@
(iv) If $S$ is reducible in $D_{n}$ while $S^{\prime}$ is not reducible in $D_{n-1}$, then the reduction must use $(i_{3},j_{3})$, which is in $D_{n}$ but not in $D_{n-1}$. If it is an NPN-reduction then it must use the newly added point $(n,0)$; in other cases the reduction can be shifted back to $S^{\prime}$. In the case of a PNP-reduction this additional point is (n1,1)(n-1,1) (when extending the staircase by a column of height zero), or can be shifted back to $S^{\prime}$ again. ∎
@@
Based on Claim 21, the incremental algorithm, sketched as Algorithm 1, generates all horizontal irreducible staircases. The PNP- and NPN-irreducibility can be checked based on Lemma 18. The last point $(i_{3},j_{3})$ is fixed, and the naïve implementation requires quadratic running time in $\operatorname{\mathrm{len}}(S)$. With some simple bookkeeping, it can be reduced to a backward scanning of the column sequence, resulting in linear running time.
0 2210 33210 332100 1111000 4332110
$0$ 3210 43210 332110 2111000 4432100
10 00000 $000$ 432100 2211000 4432110
000 10000 100000 432110 2211100 4432210
100 11000 110000 432210 2221100 5432100
110 11100 111000 433210 3221100 5432110
210 21100 211000 443210 3321100 5432210
$00$ 22100 211100 543210 3322100 5433210
1000 22110 221100 0000000 3322110 5443210
1100 32100 321100 1000000 4322100 5543210
2100 32110 322100 1100000 4322110 6543210
2110 32210 322110 1110000 4332100
Table 7. Column sequences of extremal horizontal staircases.
Using the algorithm, we have computed the complete set of irreducible staircases up to $n=60$. Table 7 lists, up to length $6$, the column sequences of those irreducible horizontal staircases that remained irreducible in each subsequent generation. The irreducible downward closed sets are these and their mirror images. Column sequences marked by a yellow background are the diagonals $D_{n}$. An example is the column sequence $2221100$ that defines the following downward closed set:
$\scriptstyle 0$$\scriptstyle 1$$\scriptstyle 2$$\scriptstyle 0$$\scriptstyle 1$$\scriptstyle 2$$\scriptstyle 3$$\scriptstyle 4$$\scriptstyle 5$$\scriptstyle 6$$\scriptstyle 1$$\scriptstyle 1$$\scriptstyle 1$$\scriptstyle 1$$\scriptstyle 1$$\scriptstyle 1$$\scriptstyle 1$$\scriptstyle 1$$\scriptstyle 2$$\scriptstyle 3$$\scriptstyle 4$$\scriptstyle 5$$\scriptstyle 1$$\scriptstyle 3$$\scriptstyle 6$$\scriptstyle 10$$\scriptstyle 15$$\scriptstyle 21$$\scriptstyle 6$$\scriptstyle 28$$\scriptstyle 7$
while the corresponding triplet is $\langle 32,76,35\rangle$. For each length, the number of sequences in Table 7 is
[…]
The number of new extremal triplets is 2x12x-1, that is,
[…]
This latter sequence matches, up to the $60$th term, the sequence A103116 in the Encyclopedia of Integer Sequences [31], which is defined as
@@
where $\varphi$ is Euler’s totient function. The formula suggests that the connection is based on the number of different slopes determined by the lattice points in a rectangle. Proving the equivalence of these two sequences is an intriguing open problem.
For better visualization, triplets $\langle\alpha_{S},\beta_{S},\gamma_{S}\rangle$ corresponding to these irreducible staircases are plotted as the three-dimensional points $\langle\beta/\alpha,\gamma/\alpha,\alpha\rangle$ using logarithmic scale for the third $\alpha$ coordinate. The plot in Figure 2 contains all 126,981 extremal triplets in the range $\beta,\gamma\leqslant 20\alpha$. Some of the plotted triplets appear as late as generation $n=80$; later generations do not contribute to this part of the complete set. For comparison, some triplets in the $80$th generation have values larger than $2^{85}$.
@@
Figure 2. Extremal configurations. Colors indicate the $\alpha$ value.
@@
$\sum_{i\leqslant x,\;j\leqslant y}\mathbfit b(i,j)$ $\approx\mathbfit b(x\mkern 1.5mu{+}\mkern 1.5mu1,y\mkern 1.5mu{+}\mkern 1.5mu1),\mbox{ and}$
$\sum_{i\leqslant x,\;j\leqslant y}i\mkern 1.5mu\mathbfit b(i,j)$ $\approx x\mkern 1.5mu\mathbfit b(x\mkern 1.5mu{+}\mkern 1.5mu1,y\mkern 1.5mu{+}\mkern 1.5mu1),$
the vector $\mathbf{v}_{S(a,b)}$ is well approximated by $\mathbfit b(x\mkern 1.5mu{+}\mkern 1.5mu1,y\mkern 1.5mu{+}\mkern 1.5mu1)\langle 1,x,y\rangle$, where $(x,y)\in S(a,b)$ is the point where $\mathbfit b(x\mkern 1.5mu{+}\mkern 1.5mu1,y\mkern 1.5mu{+}\mkern 1.5mu1)$ takes its maximal value. As the function $\mathbfit b(x,y)$ strictly increases in both coordinates, this maximum is taken on the boundary diagonal of the right-angled triangle $S(a,b)$ that has endpoints $(0,b)$ and $(a,0)$. Using the Stirling formula $n!\approx\sqrt{2\pi n}\mkern 1.5mu(n/e)^{n}$, we have
$\mathbfit b(x,y)=\frac{\mkern 1.5mu(x+y)!\mkern 1.5mu}{x!\,y!}\approx\frac{1}{\sqrt{2\pi}}\frac{(x+y)^{x+y+1/2}}{x^{x+1/2}\mkern 1.5muy^{y+1/2}}.$
Introducing $\varphi(x)=(x+\frac{1}{2})\log x$, we see that the logarithm of $\mathbfit b(x,y)$ is well approximated by the function
@@
[…]
Figure 3. The $\theta(u+1,v+1)$ function. Colors indicate the function value.